OpenStudy (anonymous):

Select the correct answer from each drop-down menu. Consider the function y=f(x)=(1/2)^x.

OpenStudy (anonymous):

The value of f(-7) is ? The value of f(5), rounded to the nearest ten thousandth, is ?

OpenStudy (anonymous):

@matthew1286

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@perl @Megatronus117

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to replace x with -7, so we get: \[f\left( { - 7} \right) = {\left( {\frac{1}{2}} \right)^{ - 7}} = ...?\]

OpenStudy (michele_laino):

now, you have to apply this property: \[{\left( {\frac{1}{a}} \right)^{ - m}} = {a^m}\] what do you get?

OpenStudy (michele_laino):

@Smileyside